OpenStudy (anonymous):

train A took 35 min to cover 50 km. if the speed of train B is 25% faster than A,it will cover distance in ?

OpenStudy (anonymous):

28 min

OpenStudy (anonymous):

how ???

OpenStudy (anonymous):

speed of A is 600/7 km/h. speed of B will be (600/7)*1.25 = 750/7 time required by B will be, (7*60*50)/750 = 28

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

2 min pls

OpenStudy (anonymous):

nope yaar !!

OpenStudy (anonymous):

well try this, In 35 min A travels = 50 km In 01 min A travels = 50/35 km In 60 min A travels = 50*60/35 km = 600/7 km Now B travels 25% faster than A. That is 1.25 times A. So, speed of B = 600*1.25/7 = 750/7 km/h Now, B travels 750/7 km in = 60 min B travels 1 km in = 60*7/750 min So, B travels 50 km in = 60*7*50/750 min 28 min. Thats ok?

OpenStudy (anonymous):

thnxx i got it, i went 4 a break so late reply -toriqur